OpenStudy (anonymous):

log8(x+1)-log8(x-4)=1.15 I solved the problem but I want to check if I did it correctly

OpenStudy (anonymous):

what did you get for x?

OpenStudy (anonymous):

I got x = -4

OpenStudy (anonymous):

ln8(x+1)-ln8(x-4) =ln8+ln(x+1)-ln8-ln(x-4) =ln(x+1)-ln(x-4) =ln[(x+1)/(x-4)]=1.15 (x+1)/(x-4)=10^1.15 (x+1)=14.13(x-4) 13.13x=5 x=5/13.13=0.38:)

OpenStudy (anonymous):

@Egirl01 sry as if x=-4 8(x+1) become -ve and then log is not defined in that region and you have to take complex that doesn't work here:)

OpenStudy (anonymous):

@Aperogalics i think by log8(x+1) he means log base 8

OpenStudy (anonymous):

yeah, that was my fault. i meant log base 8

OpenStudy (anonymous):

i should've done a better job displaying that, however, i still feel my answer is wrong

OpenStudy (anonymous):

do you mind if I show you step by step how I got the answer?

OpenStudy (anonymous):

Ohhhh all my work goes to waste:(

OpenStudy (anonymous):

haha, i'm sorry!

OpenStudy (anonymous):

ya show step by step i will not mind:)

OpenStudy (anonymous):

i got x is approx 4.5

OpenStudy (anonymous):

10.93*4+1=9.93x x=4.5:)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=log8(x%2B1)-log8(x-4)%3D1.15&t=crmtb01&f=rc if youre never sure, ask wolfram alpha wolframalpha.com

OpenStudy (anonymous):

i believe on myself only:)

OpenStudy (anonymous):

i corrected an error myself

OpenStudy (anonymous):

but i got 4.09 instead of 4.5, i'll show you what i did

OpenStudy (anonymous):

log base 8(x+1)/log base 8(x-4) 8^1.15 = x+1/x-4 (x-4)10.928 = x+1/x-4 (x-4) 10.928x – 43.713 = x+1 Subtract 10.928x from both sides -43.713 = -10.928x +1 Subtract 1 from both sides -44.713 = -10.928x Divide -10.928 on both sides X = 4.09(slightly changed, sorry, made a minor error that I fixed before. However, this answer still may be wrong)

OpenStudy (anonymous):

@Egirl01 you have done it wrong:(

OpenStudy (anonymous):

Subtract 1 from both sides -44.713 = -10.928x------------------wrong -44.713=-9.928x (as there is x already on right hand side so 1 is subtracted) :)

OpenStudy (anonymous):

that's what i thought when I saw how you got it. I made the correction and I got the same answer as you

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

done gave medals and close the questions........lol